# LLM-Enhanced Static Analysis for Precise Identification of Vulnerable OSS Versions

### v1

@@
Abstract.
Open-source software (OSS) has experienced a surge in popularity, attributed to its collaborative development model and cost-effective nature. However, the adoption of specific software versions in development projects may introduce security risks when these versions bring along vulnerabilities. Current methods of identifying vulnerable versions typically analyze and trace the code involved in vulnerability patches using static analysis with pre-defined rules. They then use syntactic-level code clone detection to identify the vulnerable versions. These methods are hindered by imprecisions due to (1) the inclusion of vulnerability-irrelevant code in the analysis and (2) the inadequacy of syntactic-level code clone detection. This paper presents Vercation, an approach designed to identify vulnerable versions of OSS written in C/C++. Vercation combines program slicing with a Large Language Model (LLM) to identify vulnerability-relevant code from vulnerability patches. It then backtraces historical commits to gather previous modifications of identified vulnerability-relevant code. We propose semantic-level code clone detection to compare the differences between pre-modification and post-modification code, thereby locating the vulnerability-introducing commit (vic) and enabling to identify the vulnerable versions between the patch commit and the vic. We curate a dataset linking 74 OSS vulnerabilities and 1013 versions to evaluate Vercation. On this dataset, our approach achieves the F1 score of 92.4%, outperforming current state-of-the-art methods. More importantly, Vercation detected 134 incorrect vulnerable OSS versions in NVD reports.
Open-source software security, Vulnerable version, Large Language Model
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Security and privacy Software security engineering
1. Introduction
@@
Public vulnerability repositories collect vulnerability reports of software products and disseminate information regarding the affected versions of the software. The National Vulnerability Database (NVD) (NVD, 2009), recognized as the largest public vulnerability database, employs the Common Platform Enumeration (CPE) format to store information about vulnerable versions. However, the NVD often encompasses all versions before the reported vulnerability or designates only the versions mentioned in the report as vulnerable. For instance, CVE-2018-5785 reported only version v2.3.0 as vulnerable in CPE (CVE, 2018). Yet, upon manual validation, it was uncovered that versions v2.1.1 to v2.3.0 are all susceptible to the vulnerability. Recent research (Dong et al., 2019; Mu et al., 2018; Tan et al., 2021; Nguyen and Massacci, 2013) indicates that incomplete and incorrect information about vulnerable versions is prevalent in such reports. Therefore, there is a need for an automated method that can more accurately identify vulnerable versions of released OSS vulnerabilities.
@@
Two limitations of current approaches. One approach for confirming vulnerabilities involves utilizing Proof of Concept (PoC) to trigger the vulnerability. PoC triggers vulnerabilities through dynamic execution, offering conclusive evidence of their existence. However, executing a PoC for each software version is a time-intensive process that requires meticulous environment setup. Additionally, the PoC input may not be universal across all vulnerable versions (Dai et al., 2021). Therefore, current research has proposed the use of static analysis to identify vulnerable versions (Da Costa et al., 2016; Rosa et al., 2021; Nguyen et al., 2016; Rodríguez-Pérez et al., 2020; Xiao et al., 2020; Shi et al., 2022; Jang et al., 2012; Woo et al., 2022). These methods typically consist of two steps: extracting vulnerability features from code snippets and identifying vulnerable versions through code clone detection of these vulnerability features. The effectiveness of vulnerability feature extraction heavily relies on human expertise in formulating vulnerability rules. Existing methods typically extract vulnerability features using pre-defined patterns: some consider data and control dependencies of the patched code (Xiao et al., 2020; Shi et al., 2022); some consider using hashes of entire functions (Woo et al., 2021); and some consider only the patch code (Jang et al., 2012; Bao et al., 2022). These methods generally produce many false positives or false negatives because the features do not adequately characterize the vulnerability (❶). Existing methods for vulnerable code clone detection are limited to syntactic signatures (such as Levenshtein algorithm (Bao et al., 2022) and hash value comparison (Xiao et al., 2020; Woo et al., 2021)), without taking into account code changes that still preserve the semantics of functions (❷).
@@
These limitations primarily stem from the lack of reasoning about the semantic context of the vulnerable code being analyzed. Recently, large language models (LLMs) have shown remarkable performance in code-related tasks including code generation (Du et al., 2024; Li et al., 2023) and code summarization (Ahmed et al., 2024; Ahmed and Devanbu, 2022). As such, we aim to address the first limitation (❶) by leveraging the power of LLM and in combination with static analysis. However, there is a challenge of Prompt engineering in leveraging a LLM in our problem domain. The design of prompts is pivotal in directing the LLM to produce desired responses. In vulnerability comprehension tasks, well-crafted prompts should encompass the code for analysis and offer a lucid description of the analysis objectives. However, in the case of current LLMs, merely providing all the vulnerable function codes and directly instructing vulnerability analysis often yields suboptimal outcomes. When the function code is too long, the LLM’s ability to understand the relationships between distant parts of the context may diminish (Li et al., 2024; Song et al., 2024). Hence, formulating prompts to guide LLMs in generating desired vulnerability logic analyses poses a challenge.
@@
Due to variations in programming styles, variable naming, and structural differences, code clones may manifest different changes. Furthermore, developers often streamline and refactor code through function outlining. Such syntactic changes often mislead syntactic code clone detection-based approaches to incorrectly identify the vulnerable versions. This motivates us to propose a semantic-level code clone detection approach to address the second limitation (❷).
@@
Our approach. We propose Vercation [^1] Vercation: Vulnerable version identification, a novel method to identify vulnerable versions of open-source C/C++ software by means of a symbiotic combination of static analysis, LLM, and semantic-level code clone detection. Given a vulnerability patch commit, Vercation applies program slicing to extract vulnerability-related statements as vulnerability features, leverages the capability of LLM in code understanding to refine the extracted features, and performs semantic-level clone detection on vulnerability features in function code changes. This hybrid method effectively overcomes the practical limitations of existing approaches. More specifically, Vercation automatically preprocess the patch commits and construct prompts based on the Few-shot and Chain-of-Thought (CoT) strategies, enabling the LLM to reason with the vulnerability and identify most probable vulnerable statements as features. Subsequently, Vercation traces earlier modifications of vulnerability features and applies a semantic clone detection method based on expanded Abstract Syntex Trees (ASTs) to pinpoint the vulnerability-introducing commit vic. Finally, Vercation identifies vulnerable versions between the vic and patch commit.
@@
Evaluation. The first author meticulously curated a ground-truth dataset for evaluation, encompassing 11 commonly used Open Source Software (OSS) projects, 74 Common Vulnerabilities and Exposures (CVEs), and a total of 1,013 OSS versions. This dataset comprised every patch released through Git with respect to those 74 CVEs. The author engaged in manual examinations, meticulously labeling the presence of vulnerabilities across an extensive set of 10,476 software-version pairs. On this dataset, Vercation demonstrated both higher precision (89.2%) and recall (95.8%) compared to state-of-the-art methods (SOTAs), including V-SZZ (Bao et al., 2022) and V0Finder (Woo et al., 2021). We conducted an ablation study using two different LLMs (GPT-4 and CodeLlama) to evaluate their vulnerability comprehension capabilities. Utilizing the Few-shot and CoT combined strategy, GPT-4 achieved an F1 score of 92.4%, significantly outperforming the F1 score of Joern parser (Joe, 2014) by 44%, a commonly-used static analysis tool. More importantly, during the evaluation, we found 134 version errors in the NVD reports. Vercation has also shown to be efficient, analyzing each vulnerability on an average of 28.61 seconds.
@@
  • We present Vercation, a novel, automated vulnerable version identification approach powered by LLM. Previous efforts heavily relied on pre-defined patterns of static analysis tools, while Vercation leverages the capability of LLM for vulnerability comprehension task, through the use of a multi-strategy universal prompt engineering.
@@
  • Vercation presents a solution based on expanded AST to address the semantic-level clone detection challenge, which was not considered in existing techniques.
@@
  • We meticulously assembled a comprehensive dataset that included 10,476 unique vulnerability-version pairs derived from 1,013 versions encompassing 74 CVEs. This extensive dataset was curated across 11 OSS projects and underwent meticulous labeling achieved through a combination of PoC input validation and rigorous manual verification.
@@
  • We have implemented a prototype of our approach and assessed its performance using our dataset, achieving the F1 score of 92.4%, outperforming SOTAs by 9% to 48%. More importantly, by applying our approach, we have detected 134 incorrect vulnerable OSS versions in NVD reports.
@@
The source code of Vercation and our curated dataset is publicly available at https://anonymous.4open.science/r/Vercation_A.
2. Background and Motivation
@@
In this section, we clarify the target problem of Vercation, introduce the LLM and discuss the motivation with two examples.
2.1. Problem Statement
@@
Commits serve as comprehensive records of OSS development, functioning as vital checkpoints for tracking the chronological evolution of code changes. They enable developers to revisit specific points on a particular date or time. This paper centers on vulnerable version identification in OSS vulnerabilities. Within the development of an OSS, certain commits introduce vulnerabilities, referred to as vulnerability-introducing commits (vic), which are later fixed in patch commits (pc). The vulnerable version analysis aims to pinpoint the vic, allowing us to assess which versions are susceptible to the vulnerabilities.
@@
We posit that the initial function undergoes modification by a vic and transforms into a vulnerable function $F_{v}$. There probably exists some subsequent commits such as feature-adding commits and refactoring commits to optimize the function code, which is denoted as $F_{vr}$. Ultimately, software analysts discover the vulnerability and patch the code in the pc, resulting in the final function $F_{p}$. An illustrative timeline is presented in Figure 1.
Figure 1. Function Timeline from Vulnerability Introduction to Vulnerability Fixing.
2.2. Large Language Model
@@
Large Language Models (LLMs) have revolutionized natural language processing and have demonstrated remarkable capabilities in various code-related tasks. These models, extensively trained on vast amounts of text and code repositories, can understand and generate human-like text and code. Recent advancements in LLMs, such as GPT-3 (Brown et al., 2020), GPT-4 (Achiam et al., 2023), and CodeLlama (Roziere et al., 2023), have demonstrated impressive performance in code generation (Du et al., 2024; Li et al., 2023), code summarization (Ahmed et al., 2024; Ahmed and Devanbu, 2022) and program repair (Fan et al., 2023; Hossain et al., 2024)..
@@
The success of LLMs in code-related tasks can be attributed to their ability to capture complex patterns and semantic relationships in code. This makes them particularly suitable for tasks that require deep understanding of code structure and functionality like vulnerability detection. However, the effectiveness of LLMs heavily depends on the quality of prompts used to guide their responses. Prompt engineering, the process of designing effective prompts, has become a crucial area of research in leveraging LLMs for specific tasks (Zhou et al., 2022; Chen et al., 2023).
@@
Despite their power, LLMs also face challenges in code analysis tasks. These include potential biases in their training data and difficulties in handling very long code sequences. In our work, we leverage the semantic understanding capabilities of LLMs to enhance the accuracy of static analysis tool in vulnerability-related code extraction, a novel application in the field of software security analysis.
2.3. Motivating Examples
@@
We present two examples of disclosed vulnerable code in two distinct scenarios, which motivate our work. Both of these vulnerabilities have now been effectively resolved by the development team. The key points we wish to underscore are as follows: 1) The significance of understanding vulnerability behavior in the process of discovering vic. 2) The challenge of accurately pinpointing vic due to the significant syntax changes introduced by a refactoring commit.
@@
2index 70142c4..012d8f5 100644
3--- a/libavformat/mxfdec.c
4$+$ b/libavformat/mxfdec.c
5@@ -493,11 +493,11 @@ static int mxf\_read\_primer\_pack
6static int mxf_read_primer_pack(void *arg, AVIOContext *pb, int tag, int size, UID uid, int64_t klv_offset)
7{

### v2

# Vercation: Precise Vulnerable Open-source Software Version Identification based on Static Analysis and LLM

@@
Abstract
Open-source software (OSS) has experienced a surge in popularity, attributed to its collaborative development model and cost-effective nature. However, the adoption of specific software versions in development projects may introduce security risks when these versions bring along vulnerabilities. Current methods of identifying vulnerable versions typically analyze and extract the code features involved in vulnerability patches using static analysis with pre-defined rules. They then use code clone detection to identify the vulnerable versions. These methods are hindered by imprecision due to (1) the exclusion of vulnerability-irrelevant code in the analysis and (2) the inadequacy of code clone detection. This paper presents Vercation, an approach designed to identify vulnerable versions of OSS written in C/C++. Vercation combines program slicing with a Large Language Model (LLM) to identify vulnerability-relevant code from vulnerability patches. It then backtracks historical commits to gather previous modifications of identified vulnerability-relevant code. We propose code clone detection based on expanded and normalized ASTs to compare the differences between pre-modification and post-modification code, thereby locating the vulnerability-introducing commit (vic) and enabling the identification of the vulnerable versions between the vulnerability-fixing commit and the vic. We curate a dataset linking 122 OSS vulnerabilities and 1,211 versions to evaluate Vercation. On this dataset, our approach achieves an F1 score of 93.1%, outperforming current state-of-the-art methods. More importantly, Vercation detected 202 incorrect vulnerable OSS versions in NVD reports.
Index Terms:
Open-source software security, Vulnerable version, Large Language Model.
footnotetext: 5 Shichao Lv and Ting Zhang are the corresponding authors.
I Introduction
@@
Public vulnerability repositories collect vulnerability reports of software products and disseminate information regarding the affected versions of the software. The National Vulnerability Database (NVD) [1], recognized as the largest public vulnerability database, employs the Common Platform Enumeration (CPE) format to store information about vulnerable versions. However, the NVD often encompasses all versions before the reported vulnerability or designates only the versions mentioned in the report as vulnerable. For instance, CVE-2018-5785 reported only version v2.3.0 as vulnerable in CPE [2]. Yet, upon manual validation, it was uncovered that versions v2.1.1 to v2.3.0 are all susceptible to the vulnerability. Recent research [3, 4, 5, 6] indicates that incomplete and incorrect information about vulnerable versions is prevalent in such reports. Therefore, there is a need for an automated method that can more accurately identify vulnerable versions of released OSS vulnerabilities.
@@
Two limitations of current approaches. The approaches for confirming vulnerabilities involve utilizing Proof of Concept (PoC) to trigger the vulnerability. PoC triggers vulnerabilities through dynamic execution, offering conclusive evidence of their existence. However, executing a PoC for each software version is a time-intensive process that requires meticulous environment setup. Additionally, the PoC input may not be universal across all vulnerable versions [7]. Therefore, current researches use static analysis to identify vulnerable versions [8, 9, 10, 11, 12, 13, 14, 15]. These methods typically consist of two steps: extracting vulnerability features from code snippets and identifying vulnerable versions through code clone detection of these vulnerability features. Existing methods typically extract vulnerability features using static analysis in which vulnerability patterns are pre-defined by human experts: some consider data and control dependencies from the patched code [12, 13]; some consider using hashes of entire functions [16]; and some consider only the patch code [14, 17]. These methods generally produce many false positives or false negatives because of the inherently imprecise nature of static analysis and the incompleteness of pre-defined vulnerability patterns, making it difficult to distinguish vulnerability-relevant code from irrelevant code (❶). The majority of the methods for code clone detection are function-level approaches [18] whereas real-world vulnerability logic is typically confined to (often a few) statements only. There are statement-level code clone detection methods but they are limited to textual signatures (such as Levenshtein algorithm [17] and hash value comparison [12, 16]), without taking into account the code changes that reorganize code while preserving its functionality. These limitations hinder the effectiveness of vulnerability detection in practical scenarios (❷).
@@
These limitations primarily stem from the lack of reasoning about the context of the vulnerable code being analyzed. Recently, large language models (LLMs) have shown remarkable performance in various code-related tasks, including code generation [19, 20] and code summarization [21, 22]. This strong performance demonstrates LLMs’ capability to recognize code patterns and correlations at both syntactic and semantic levels. As such, we aim to address the first limitation by leveraging the power of LLMs. Essentially, we hypothesize that LLMs can be leveraged to address the imprecision of static analysis and the incompleteness of human-defined vulnerability patterns (❶). However, there is a challenge of Prompt engineering in leveraging an LLM in our problem domain. The design of prompts is pivotal in directing the LLM to produce desired responses. In vulnerability comprehension tasks, well-crafted prompts should encompass the code for analysis and offer a lucid description of the analysis objectives. However, in the case of current LLMs, merely providing all the vulnerable function codes and directly instructing vulnerability analysis often yields suboptimal outcomes. When the function code is too long, the LLM’s ability to understand the relationships between distant parts of the context may diminish [23, 24]. Hence, formulating prompts to guide LLMs in generating desired vulnerability logic analyses poses a challenge.
@@
Due to variations in programming styles, variable naming, and structural differences, code clones may manifest different changes. Furthermore, developers often streamline and refactor code through method outlining. Such code structural changes often mislead code clone detection-based approaches to incorrectly identify the vulnerable versions. Furthermore, prior studies [25, 26] found that vulnerabilities can often be localized to a few key lines, making file- and function-level vulnerability detection overly coarse-grained. The limitation motivated us to propose a statement-level code clone detection approach to address the second limitation of current approaches (❷).
@@
Our approach. We propose Vercation (Vulnerable version identification), a novel method to identify vulnerable versions of open-source C/C++ software utilizing a symbiotic combination of static analysis, LLM, and code clone detection. Given a vulnerability fixing commit (vfc), Vercation applies program slicing to extract vulnerability-related statements as vulnerability features, leverages the capability of LLM in code understanding to refine the extracted features, and performs semantic-level clone detection on vulnerability features in code changes. This hybrid method effectively overcomes the practical limitations of existing approaches. More specifically, Vercation automatically preprocess the fixing commits and construct prompts based on the Few-shot and Chain-of-Thought (CoT) strategies, enabling the LLM to reason with the vulnerability and identify the most probable vulnerable statements as features. Subsequently, Vercation traces earlier modifications of vulnerability features and applies a clone detection method based on expanded Abstract Syntax Trees (ASTs) to pinpoint the vulnerability-introducing commit vic. Finally, Vercation identifies vulnerable versions between the vic and vfc.
@@
Evaluation. We meticulously curated a ground-truth dataset for evaluation, encompassing 12 commonly used OSS projects, 122 Common Vulnerabilities and Exposures (CVEs), and a total of 1,211 OSS versions. This dataset comprised every patch released through Git with respect to those 122 CVEs. The first author engaged in manual vulnerability validation with the help of public Proof-of-Concept (PoC), meticulously labeling the presence of vulnerabilities across software versions. On this dataset, Vercation demonstrated both higher precision (91.8%) and recall (94.5%) compared to state-of-the-art methods (SOTAs), including V-SZZ [17], V0Finder [16], V1SCAN [27], VERJava [28] and Vision [29]. We conducted an ablation study using three different LLMs (GPT-4 [30], CodeLlama [31], and DeepSeek-V3 [32]) to evaluate their vulnerability comprehension capabilities. Utilizing the Few-shot and CoT combined strategy, DeepSeek-V3 achieved an F1 score of 93.1%, significantly improving the F1 score of Joern parser [33] by 92.8%, a commonly-used static analysis tool. More importantly, during the evaluation, we found 202 version errors in the NVD reports. Vercation has also been shown to be efficient, analyzing each vulnerability on an average of 28.61 seconds.
@@
  • Unlike previous efforts that heavily relied on pre-defined patterns of static analysis tools, we present Vercation, the first framework to integrate the reasoning capability of LLM for vulnerable version identification tasks, through the use of a multi-strategy universal prompt engineering.
@@
  • Vercation presents a solution based on expanded and normalized AST to address the structural modifications in the clone detection challenge, which was designed for refactoring commit identification during vulnerability backtrack.
@@
  • We curated a dataset including 122 published CVEs containing 1,211 versions. This extensive dataset was curated across 12 OSS projects and underwent meticulous labeling through a combination of PoC input validation and manual verification.
@@
  • We have implemented a prototype of our approach and assessed its performance using our dataset, achieving the F1 score of 93.1%, improving SOTAs by 8.1% to 108.7%. We also evaluated three contrasting LLMs — a commercial, closed model (GPT-4) and open models (CodeLlama and DeepSeek-V3) — in our approach and reported that they achieved similar performances in our context when appropriate prompting strategies were applied. More importantly, by applying our approach, we have detected 202 incorrect vulnerable OSS versions in NVD reports.
@@
The source code of Vercation and our curated dataset are publicly available at https://github.com/Veronica-L/Vercation.
II Background and Motivation
@@
In this section, we clarify the target problem, introduce the LLM and discuss the motivation with two examples.
II-A Problem Statement
@@
Commits serve as comprehensive records of OSS development, functioning as vital checkpoints for tracking the chronological evolution of code changes. They enable developers to revisit specific points on a particular date or time. This paper centers on vulnerable version identification in OSS vulnerabilities. Within the development of an OSS, certain commits introduce vulnerabilities, referred to as vulnerability-introducing commits (vic), which are later fixed in vulnerability fixing commits (vfc). The vulnerable version analysis aims to pinpoint the vic, allowing us to assess which versions are susceptible to the vulnerabilities.
@@
We posit that the initial function undergoes modification by a vic and transforms into a vulnerable function $F_{v}$. There probably exists some subsequent commits such as feature-adding commits and refactoring commits to optimize the function code, which is denoted as $F_{vr}$. Ultimately, software analysts discover the vulnerability and fix the code in the vfc, resulting in the final function $F_{f}$. An illustrative timeline is presented in Figure 1.
Figure 1: Function Timeline from Vulnerability Introduction to Vulnerability Fixing.
II-B Large Language Model
@@
Large Language Models (LLMs) have revolutionized natural language processing and have demonstrated remarkable capabilities in various code-related tasks. These models, extensively trained on vast amounts of text and code repositories, can understand and generate human-like text and code. Recent advancements in LLMs, such as GPT-3 [34], GPT-4 [30], DeepSeek [35] and CodeLlama [31], have demonstrated impressive performance in code generation [19, 20], code summarization [21, 22] and program repair [36, 37].
@@
The success of LLMs in code-related tasks can be attributed to their ability to capture complex patterns and contextual relationships in code [38, 31]. Through Pattern Recognition, LLMs learn correlations and patterns between code snippets from large-scale training data, enabling them to identify code with different syntax but similar functionality. Additionally, LLMs can infer the purpose of code snippets based on their context. This makes them particularly suitable for tasks that require a deep understanding of code structure and functionality like vulnerability detection. However, the effectiveness of LLMs heavily depends on the quality of prompts used to guide their responses. Prompt engineering, the process of designing effective prompts, has become a crucial area of research in leveraging LLMs for specific tasks [39, 40].
@@
Despite their reasoning capabilities, LLMs also face challenges in code analysis tasks such as handling very long code sequences. Therefore, in our work, instead of blindly applying LLM as a standalone tool, we first use static analysis to extract candidate vulnerability-related codes and then leverage the semantic capturing capabilities of LLMs to improve the accuracy of extracted codes.
II-C Motivating Examples
@@
We present two examples of disclosed vulnerable code in two distinct scenarios, which motivate our work. Both of these vulnerabilities have now been effectively resolved by the development team. The key points we wish to underscore are as follows: 1) The significance of understanding vulnerability behavior in the process of discovering vic. 2) The challenge of accurately pinpointing vic due to the code structural modifications introduced by the intermediate commits.
Example 1) A vulnerability in FFmpeg (CVE-2017-14169 [41]) shows the limitation of existing approaches in terms of capturing vulnerability logic. The code snippet of vfc is shown in Listing LABEL:list:CVE-2017-14169. A sanitizing check was added at Line 13 for the item_num. Without this check, a remote attacker can make a crafted file with a large item_num field such as 0xffffffff, causing a buffer overflow issue in the avio_read function at Line 25 and potentially leads the application to exhibit incorrect behavior or crash.
V-SZZ [17] assumes the deletion lines in the security patch as vulnerable codes and the basic idea of it is to pinpoint the earliest commit introducing the deletion lines as vic. However, vulnerability logic is composed of various vulnerable codes, such as Lines 21 and 25 in Listing LABEL:list:CVE-2017-14169. V-SZZ only considers tracing back deleted line (Line 12) and overlooks the real vulnerability behavior, leading to incorrect identification of vic. V0Finder [16] attempts to identify vulnerable versions using code clone detection. It generates hash values for the entire vulnerability function and identifies code clones according to the distance value of two hashes. This method may still introduce excessive vulnerability-unrelated features due to the inclusion of “entire function”.
@@
2@@ -493,11 +493,11 @@ static int mxf\_read\_primer\_pack
3static int mxf_read_primer_pack(void *arg, AVIOContext *pb, int tag, int size, UID uid, int64_t klv_offset)
4{